# Unleashing Expert Opinion from Social Media for Stock Prediction

### v1

@@ VI-C Ablation Study
In this ablation study, we analyze the impact of different graph structures and graph neural network architectures on the performance of our model. For the graph structure, we tested three configurations using the GAT architecture: employing only the industry graph, only the correlation graph, and a combination of both by our DualGAT. The results shown in Table III indicated that the combined graph setup outperformed the individual setups, demonstrating the effectiveness of integrating diverse market relationships into the model. Meanwhile, for a single graph source, we find that combining the information contained in the pre-trained model output with the expert signal yields better results than the message passing using only expert signals, especially in terms of AR and SR. This further demonstrates that expert signals can work synergistically with traditional financial features.
@@ VII Discussion
A critical aspect of our framework is its robustness in handling sparse expert signals across different sectors and stock types. Our dual-graph architecture effectively addresses the sparsity problem through different propagation mechanisms. We have found that the industry graph enables each stock’s signal to reach an expected 20.65% of nodes in a single hop while the correlation graph complements this by propagating signals to an additional 9.16% of nodes. This dual-graph approach ensures comprehensive and sufficient signal coverage around 90% after two-hop propagation. The accuracy of signal propagation is equally crucial as coverage. Our analysis reveals strong trend alignment between connected stocks, supporting the reliability of our propagation mechanism. In the correlation graph, stock pairs meeting our correlation thresholds (0.67 and 0.77) show trend alignment probabilities of 75.09% and 80.04% respectively in the next day. Similarly, stocks connected in the industry graph demonstrate a 70.4% probability of aligned trends. Even accounting for potential error propagation, the accuracy of the propagated expert signals remains robust, with even the simplest message passing approach could outperform baseline models in Table II.
@@
Ultimately, our method emphasizes the use of expert opinions based on self-labeled sentiments, avoiding the computational complexity of directly processing large volumes of tweets using language models. Though large language models (LLMs) have become increasingly mature and widely used in recent years, the massive volume of posts on StockTwits poses significant computational challenges. Processing over 200,000 daily tweets with LLMs would significantly slow down inference times, result in missing many critical information thus make real-time predictions impractical for our task setting since decisions must be made before the market close. As such, our approach does not compete with LLMs but rather complements them by focusing on meaningful social media opinions. For instance, one promising future direction involves using our expert tracing system to quickly identify potential experts in the large volume of daily tweets. This would enable us to prioritize their texts for LLM processing, allowing the LLM to extract additional information beyond sentiment. Such selective processing would enhance the synergy between the capability of LLMs and expert signals.
@@ VIII Conclusion
For future work, there is substantial potential to further exploit the wealth of information embedded in social media content. As StockTwits dataset also contains the follower count for each user, we could investigate the relationship between social media influence and stock price movements, as well as examine the predictive capabilities of influential users. Another promising direction could involve processing the textual content of tweets by the experts or social media influencers through large language models to generate embeddings. These embeddings could then be integrated with sentiment features, providing a richer set of features for our graph-based models.

### v2

@@ VI-C Ablation Study
In this ablation study, we analyze the impact of different components on the performance of our model.
For the graph structure, we tested three configurations using the GAT architecture: employing only the industry graph, only the correlation graph, and a combination of both by our DualGAT. The results shown in Table III indicated that the combined graph setup outperformed the individual setups, demonstrating the effectiveness of integrating diverse market relationships into the model. Meanwhile, for a single graph source, we find that combining the information contained in the pre-trained model output with the expert signals yields better results than the message passing using only expert signals, especially in terms of AR and SR. This further demonstrates that expert signals can work synergistically with traditional financial features.
@@
Beyond the architectural components, we also evaluated the impact of expert signal quality on our framework. We fed our DualGAT model with expert signals generated by various baseline expert identification algorithms shown in Figure 1. Our proposed algorithm consistently outperforms all baseline methods, confirming its critical contribution to the model’s superior performance. Detailed results are available in Supplementary Material S9.
@@ VII Discussion
A critical aspect of our framework is its robustness in handling sparse expert signals across different sectors and stock types. Our dual-graph architecture effectively addresses the sparsity problem through different propagation mechanisms. We have found that the industry graph enables each stock’s signal to reach an expected 20.65% of nodes in a single hop while the correlation graph complements this by propagating signals to an additional 9.16% of nodes. This dual-graph approach ensures comprehensive and sufficient signal coverage around 90% after two-hop propagation. Empirical analysis further confirms that two-hop propagation achieves optimal performance (see Supplementary Material S10). The accuracy of signal propagation is equally crucial as coverage. Our analysis reveals strong trend alignment between connected stocks, supporting the reliability of our propagation mechanism. In the correlation graph, stock pairs meeting our correlation thresholds (0.67 and 0.77) show trend alignment probabilities of 75.09% and 80.04% respectively in the next day. Similarly, stocks connected in the industry graph demonstrate a 70.4% probability of aligned trends. Even accounting for potential error propagation, the accuracy of the propagated expert signals remains robust, with even the simplest message passing approach outperforming baseline models in Table II.
@@
Ultimately, our method emphasizes the use of expert opinions based on self-labeled sentiments, which avoids the computational overhead associated with large language models (LLMs). While LLMs could be applied for sentiment analysis, our approach is demonstrably more computationally efficient and better suited for real-time applications. In fact, our entire framework, from signal extraction to the DualGAT architecture, is designed for high efficiency. Empirical analysis shows that our DualGAT model’s inference time exhibits a growth rate significantly below linear scaling with the number of stocks, ensuring its practicality for large-scale use. A detailed computational and scalability analysis is provided in Supplementary Material S11. As such, our approach does not compete with LLMs but complements them by focusing on meaningful opinions. For instance, one promising direction involves using our expert tracing system to quickly identify potential experts in daily tweets. This enables prioritizing their texts for LLM processing, allowing LLMs to extract information beyond sentiment. This selective processing enhances synergy between LLM capabilities and expert signals.
@@ VIII Conclusion
For future work, there is substantial potential to further exploit the wealth of information in social media content. As StockTwits dataset contains follower counts for each user, we could investigate the relationship between social media influence and stock price movements, and examine the predictive capabilities of influential users. Another promising direction involves processing textual content of expert or influencer tweets through large language models to generate embeddings. These embeddings could be integrated with sentiment features, providing richer features for our graph-based models.